OpenStudy (he66666):

Solving linear trig equations? Solve tan(3x)=1 if -pi≤x≤pi. answer: x=pi/12, 5pi/12, 3pi/4, -pi/4, -7pi/12, -11pi/12 I don't get how you do this question.. I know that the restrictions become -3pi≤3x≤3pi but i can't do beyond that.. can somebody please help?

OpenStudy (anonymous):

Principle angle = arctan(1) = pi/4. => 3x = pi/4 Also, tan(a) = tan(b) <=> a = pi*n + b => 3x also equals pi/4 +/- pi within the restriction +/-3pi. Then divide all these by 3.

OpenStudy (he66666):

I still don't get it... I can't until 3x = pi/4 :S

OpenStudy (anonymous):

Oh dear :(. You need to go re-learn it then

OpenStudy (he66666):

I just don't get the restriction part of it.. if it's 0≤x≤2pi, I can do most questions but if it's -pi≤x≤pi or other "complicated" ones, I can't do it

OpenStudy (anonymous):

It's exactly the same principle (especially for tan) ...